OpenStudy (anonymous):

If 5.78 g of CuNO3 is dissolved in water to make a 0.580 M solution, what is the volume of the solution?

OpenStudy (anonymous):

:)

OpenStudy (rushwr):

First we will find the no. of moles \[no. of moles = \frac{ mass }{ molar mass } \] Mass is given here as 5.78g Molar mass of CuNO3 = 125.5gmol^-1 Since we know these we can use the above equation to find the moles: \[moles = \frac{ 5.78 }{ 125.5 } = 0.046 moles\] There's another equation; \[Volume = \frac{ moles }{ concentration } = \frac{ 0.046 }{ 0.580 } = 0.0793dm ^{3}\]

OpenStudy (anonymous):

OKAY! okay I was .006 off with the moles thats why!

OpenStudy (rushwr):

Alright ! :)